AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 575     Entry time: Sun Jun 2 13:04:41 2024
Author: JM, AG, CB, DR 
Subject: Run 23 - vertical beam position and beam lifetime tests 

14:05

Run 22 stopped. Run 23 started we will now adjust the vertical beam position to see if we observe the beam stripe seen in the last beamtime.

Current vertical beam position 0.0 mm. Horizontal beam position 7.0 mm. Coasting beam.

Beam D+, 1.5 MeV. Storage time ~ 20 minutes. Target ON. Density ~1e11 at/cm2.

XY plot from run 22 with the same settings - attachment 1

Beam stripe looks like it is just on bottom detectors (just )

14:10

New beam settings

Vertical Beam position 5.0mm, Horizontal unchanged. Storage time reduced to ~2 minute

No real change observed. Perhaps some thing more can be seen on top right detector but definitely not like in the Feb experiment.

14:15

New beam settings

Vertical Beam position 9.0mm, Horizontal unchanged.

No real change observed again.

14:20

New beam settings

Vertical Beam position -9.0mm, Horizontal unchanged.

More events observed on bottom left detector but nothing like the Feb measurements. I think we do not have stripe effects like previously.

14:40:50

New Beam settings to check beam lifetime.

Vertical beam position set back to 0.0 mm. Horizontal unchanged

In bunched mode to observed intensity over time. 

Injection into ring 2.2e7

15:15

Beam intensity dropped to 1.1e7. Beam lifetime 35 minutes.

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

 

15:24

We removed the target. Beam intensity at 2.2E7

After 35 minutes, the beam intensity is 1.41e7. So target does have effect on beam lifetime! We would like more.

16:05

We will now bring target back in and try to increase density.

16:10

Some sort of GSI-wide failure, we have lost beam entirely. We don't see any effect on our system. Temperatures and leakage currents are stable

Detectors moved fully out for safety

17:15

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

19:15

Still no beam

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

21:15

Still no beam

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

21:40

Beam has returned!

22:00

Target at around 9e11 and appears stable

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

22:33

Beam back in the ring. Beam settings to be updated.

Rates in right hand detectors look slightly different to before - lower in bottom right, slightly higher in top left. Perhaps beam vertical position needs to be lowered slightly.

ASIC check load just to be sure. Rates in aida15 greatly reduced - only temporarily.

 

Attachment 1: Screenshot_from_2024-06-02_14-04-55.png  697 kB  Uploaded Sun Jun 2 14:11:33 2024  | Show | Show all
ELOG V3.1.4-unknown